A tennis ball rolls 10 meters east when it hits a stone that turns the ball 80 ᵒ N of E

rolling for another 15 m. Calculate magnitude and direction of the resultant displacement of the 

tennis ball.


1
Expert's answer
2021-10-13T15:40:48-0400

"x=10+15\cdot\cos80°=12.6\ (m)"

"y=80\cdot\sin80°=78.8\ (m)"


"l=\sqrt{12.6^2+78.8^2}=79.8\ (m)" . Answer